# Fibonacci-Modulation-Induced Multiple Topological Anderson Insulators

### v1

@@ Abstract
We uncover the emergence of multiple topological Anderson insulators (TAIs) in a 1D spin-orbit coupled (SOC) chain driven by Fibonacci modulation, transforming a trivial band structure into a cascade of topologically nontrivial phases. This intriguing phenomenon is marked by the appearance of zero-energy modes and transitions in the $\mathcal{Z}_{2}$ topological quantum number. Strikingly, as the SOC amplitude decreases, the number of TAI phases grows, a behavior intricately linked to the fractal structure of the energy spectrum induced by Fibonacci modulation. Unlike conventional TAI phases, which exhibit fully localized eigenstates, the wave functions in the Fibonacci-modulated TAI phases exhibit multifractal behavior. Furthermore, this model can be experimentally realized in a Bose-Einstein condensate along the momentum lattice, where its topological transitions and multifractal properties can be probed through quench dynamics. Our findings open new avenues for exploring exotic disorder-induced topological phases and their intricate multifractal nature.
Introduction.—Topological states of matter have been extensively studied in various systems, including condensed-matter materials[1, 2, 3, 4, 5], ultracold atoms[6, 7], superconducting circuits[8, 9, 10, 11], electronic circuits[12], photonic lattices[13, 14, 15, 16, 17], and mechanical systems[18]. Topological insulators are characterized by gapless edge states that are immune to backscattering from weak disorders[19, 20, 21], but these states typically cannot survive strong disorders due to Anderson localization[22, 23]. Interestingly, a reverse transition has been observed: disorder can induce the emergence of protected edge states and quantized transport even in a trivial band structure. This disorder-induced topological phase, known as the topological Anderson insulator (TAI)[24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], has been theoretically predicted in systems such as the two-dimensional (2D) HgTe/CdTe quantum well[24, 25, 35] and the 1D Su-Schrieffer-Heeger chain[36], and experimentally realized in a variety of engineered systems, including 1D cold atomic wires[37], 2D photonic waveguide arrays[38, 39], and photonic quantum walk[40]. Notably, both uncorrelated random[37, 41, 42, 43, 44, 45, 46] and quasiperiodic disorders in the form of a cosine modulation incommensurate with lattice spacing[47, 48, 49, 34, 50, 51] can give rise to the TAI phase. Expanding the range of disorder types that can induce the TAI phase, and exploring the associated exotic topological and localization effects, remains a critical focus of ongoing research.
Quasicrystals offer intriguing platforms for studying exotic localization phenomena and topologically nontrivial behavior in physics[55, 54, 53, 52, 56, 57, 58, 59]. Unlike randomly disordered systems, quasicrystals exhibit a structure that is long-range ordered but not periodic[60, 61]. Significant advancements in quasicrystal growth and synthetic engineering, achieved with atomic and photonic precision, have made it possible to create quasiperiodic structures. A notable example is the Fibonacci quasicrystal, which hosts critical states that are neither fully localized nor extended[62, 63, 64, 65, 66, 67, 61], along with a multifractal spectrum of topological gaps and subgap winding states[56, 57, 58, 59]. Synthetic Fibonacci chains have been realized in various contexts, including cold atoms[68, 69], photonics[64, 71, 70, 72], polaritonics[73, 74], and dielectric chains or circuits[75, 76], where the wave functions and topological properties of Fibonacci chains have been experimentally measured[77, 56, 78, 79, 80].
@@
In this Letter, we report the emergence of multiple topological Anderson insulators (TAIs) induced by Fibonacci modulation in a 1D spin-orbit coupled (SOC) chain, starting from a trivial band structure. This intriguing phenomenon is confirmed by the appearance of nontrivial zero modes, along with changes in the $\mathcal{Z}_{2}$ topological quantum number. The number of TAI phases increases as the SOC amplitude decreases. The fractal structure of the spectrum plays a crucial role in governing the emergence of these multiple TAI phases. Specifically, the self-similar splitting of the spectrum induced by the Fibonacci modulation facilitates the formation of multiple topologically nontrivial regions, which evolve as the SOC amplitude is varied. In contrast to conventional TAI phases, which exhibit fully localized eigenstates, the wave functions in the Fibonacci-modulated TAI phases display multifractal characteristics. Furthermore, our model can be experimentally realized in a Bose-Einstein condensate along the momentum lattice, where its topological transitions and multifractal behavior can be probed through quench dynamics.
@@
Model.—We study a 1D SOC atomic chain with a Fibonacci-modulated on-site potential, described by the Hamiltonian
@@
where $\Psi_{i}^{\dagger}=(c^{\dagger}_{i\uparrow},c^{\dagger}_{i\downarrow})$ with $c^{\dagger}_{i\beta}$ creating a particle with spin $\beta=\{\uparrow,\downarrow\}$ at site $i$. The hopping matrix takes the form $\mathcal{R}_{ij}=-t_{0}\sigma_{z}\pm it_{so}\sigma_{y}$, along the ±$\pm\hat{x}$ direction, respectively. The diagonal terms of $\mathcal{R}$ describe the spin-conserved hopping with the amplitude $t_{0}$ (set to unit, $t_{0}=1$) and its off-diagonal terms represent the spin-flip hopping between the nearest neighbors with the amplitude $t_{so}$. The second term in the Hamiltonian corresponds to an on-site Fibnacci modulation $\mathcal{U}_{i}=\left[\lambda\mathrm{sgn}\left(\cos(2\pi\alpha i+\phi)-\cos(\pi\alpha)\right)+M\right]\sigma_{z}$ [61], where $M$ denotes a constant uniform Zeeman potential, $\lambda$ is the amplitude of quasiperiodic modulation, $\textup{sgn}[\cdots]$ denotes the sign function, and $\phi\in[0,2\pi)$ is an arbitrary phase. The modulation frequency $\alpha=(\sqrt{5}-1)/2$ is derived from the Fibonacci sequence $F_{\nu+1}=F_{\nu-1}+F_{\nu}$, with $F_{0}=F_{1}=1$[81]. We take the rational approximation $\alpha=F_{\nu-1}/F_{\nu}$ with the system size chosen as $L=F_{\nu}$.
@@
The Hamiltonian $H$ respecting the chiral symmetry defined by $\sigma_{x}$ in spin space governs a Fibonacci-modulated SOC chain, whose counterpart with a constant uniform Zeeman potential can rise to the 1D AIII class topological insulator[82]. In the crystalline case, where $\mathcal{U}_{i}=M\sigma_{z}$, the clean SOC model splits into two bands with the eigenvalues $E_{\pm}=\pm\sqrt{\left(M-2\cos k\right)^{2}+4t_{so}^{2}\sin^{2}k}$, where $k$ is the wave vector. For $|M|>2$, regardless of whether $t_{so}\neq 0$, a topological phase transition occurs, marked by a gap closing at $\left|M\right|=2$. This transition is also characterized by the vanishing of zero-energy edge modes and the disappearance of nontrivial topological invariants. The main focus of this Letter is on the topological phase transition and localization effects induced by the Fibonacci modulation.
Figure 1: (Color online) (a) The disorder-averaged topological phase diagram of the system in $M-\lambda$ plane with $t_{so}=0.5$. The blue dashed lines correspond to the topological phase boundaries determined by Eq.(2). Two disorder-averaged energies $\overline{E}_{L}$ and $\overline{E}_{L+1}$ at the center of the spectrum and the disorder-averaged topological number $\overline{Q}$ as a function of $\lambda$ under OBC with $t_{so}=0.5$ for (b) $M=3$ and (c) $M=-4$. (d) The number of times the TAI phases emerge, denoted as $\overline{g}$, as a function of $t_{so}$ with $M=-4$. Here, all the data are averaged by $N_{c}=50$ disorder realizations.
@@
TAI and multiple TAIs—Topologically protected edge states and quantized topological charges emerge in the context of TAI when sufficient disorder or incommensurate modulation is introduced to a trivial band structure. We employ the scattering-matrix method to compute the $\mathcal{Z}_{2}$ topological quantum number $Q$ in our case [83], where $Q=1$ ($Q=-1$) corresponds to a topologically trivial (nontrival) phase. Figure 1(a) shows the disorder-averaged topological phase diagram with $t_{so}=0.5$ as a function of $M$ and $\lambda$, obtained numerically averaging the disorder-averaged $\mathcal{Z}_{2}$ number $\overline{Q}$ over $N_{c}$ disorder realizations for different $\phi\in[0,2\pi)$. The phase boundaries can be determined by the condition:
@@
where $T$ is the total transfer matrix defined in a new basis $\tilde{\Psi}_{i}^{\dagger}=(c_{i\uparrow}^{\dagger}+c_{i\downarrow}^{\dagger},c_{i\uparrow}^{\dagger}-c_{i\downarrow}^{\dagger})$ for the zero modes [83]. In the absence of $\lambda$, the system exhibits a topological phase transition at $|M|=2$, corresponding to a jump in $\overline{Q}$ from $-1$ to $1$ as $|M|$ increases. When the Fibonacci modulation is introduced to the nontrivial band structure with $|M|<2$, the topologically nontrivial phase remains robust against weak $\lambda$, but eventually transitions to a trivial one for sufficiently strong modulation. For $M>2$, the system undergoes a topological phase transition from trivial to nontrivial to trivial as $\lambda$ increases, indicating the emergence of the TAI. We further present two disorder-averaged energies, $\overline{E}_{L}$ and $\overline{E}_{L+1}$, at the center of the spectrum, as well as $\overline{Q}$ as a function of $\lambda$ under open boundary condition (OBC) with $M=3$ and $t_{so}=0.5$ in Fig. 1(b). When $\lambda$ lies in the range $[2.51,2.94]$, the topological number $\overline{Q}$ jumps from $1$ to $-1$, accompanied by the emergence of the zero edge modes. For $\lambda>2.94$, the topologically nontrivial phase vanishes. Our numerical results suggest that the Fibonacci modulation can induce the TAI.
@@
As seen in Fig. 1(a) for $M<-2$, the system undergoes multiple distinct transitions into the TAI phase, eventually returning to a trivial phase as $\lambda$ increases further. Figure 1(c) shows the disorder-averaged energies of the two center modes, $\overline{E}_{L}$ and $\overline{E}_{L+1}$, as well as $\overline{Q}$ as a function of $\lambda$ with $M=-4$ and $t_{so}=0.5$. It can be observed that, as the $\mathcal{Z}_{2}$ number $\overline{Q}$ transitions to a nontrivial value, corresponding zero modes emerge in the ranges $\lambda\in(2.87,3.22)\cup(3.67,3.94)\cup(4.66,4.88)$. The system undergoes three transitions from topologically trivial phases to nontrivial ones, indicating the emergence of the multiple TAI phases. For a given $M<-2$, the number of times TAI phases emerge, denoted by $g$, strongly depend on $t_{so}$. As shown in Fig. 1(d) with $M=-4$, $\overline{g}$ gradually increases as $t_{so}$ decreases, for $N_{c}=50$ disorder realizations.
@@
To uncover the underlying mechanism of this phenomenon, we analyze the energy spectrum of two decoupled Fibonacci-modulated chains without the SOC term ($t_{so}=0$) as described in Eq. (1), shown in Fig. 2(a) for $\phi=0$. For finite $\lambda$, the spectrum exhibits eight primary bands. Within the range $\lambda\in(2.76,4.95)$, the middle six bands overlap, forming three clusters of band-crossing regions separated by two prominent band gaps near zero energy. Remarkably, each cluster undergoes successive splitting into smaller sub-clusters, creating a self-similar fractal structure that persists in the thermodynamic limit. This fractal nature manifests in the emergence of progressively smaller band gaps $\Delta_{g}$ at each level of splitting. When the SOC strength $t_{so}$ is introduced, these band gaps—comparable in size to $t_{so}$—begin to close and reopen, signaling the formation of nontrivial zero-energy modes, as illustrated in Fig. 2(b). This process unveils an infinite hierarchy of TAI phases in the small $t_{so}$ regime and in the thermodynamic limit, effectively creating a topologically fractal structure. As $t_{so}$ increases further, neighboring topologically nontrivial regions merge, reducing the number of distinct TAI phases. This transition, depicted in Figs. 2(c) and 2(d), signifies a transformation from a fractal topology to a more discrete structure, highlighting how the interplay between SOC and Fibonacci modulation reshapes the landscape of topological phases.
We investigate an off-diagonal Fibonacci-modulated SOC chain and uncover the emergence of multiple TAI phases, whose occurrence strongly depends on $t_{so}$. Notably, the localization properties exhibit multifractal behavior, regardless of whether the system resides in a TAI phase or not, as detailed in the Supplementary Material [83]. These findings demonstrate that both diagonal and off-diagonal Fibonacci modulation in SOC chains can induce multiple TAI phases, offering new insights into quasiperiodic topological systems.
Figure 2: (Color online) (a) Self-similarities of energy spectrum as a function of $\lambda$ with $t_{so}=0$. Two central energy levels $E_{L}$ and $E_{L+1}$ as a function of $\lambda$ with (b) $t_{so}=0.25$, (c) $0.5$, and (d) $0.9$. Here, $M=-4$, $\phi=0$, and $L=610$.
Figure 3: (Color online) Density distributions of the selected states in both real (the left column) and momentum (the right column) spaces. The top row corresponds to the $198$th eigenstates with $\lambda=2$. The bottom row corresponds to the $198$th eigenstates with $\lambda=3$. Here, $t_{so}=0.5$, $M=-4$, $\phi=0$, and $L=610$.
Figure 4: (Color online) Fractal dimensions $D_{n}$ for different system sizes under PBC for (a) $\lambda=2$ and (b) $\lambda=3$, respectively. (c) The scaling of the mean fractal dimensions $\langle D\rangle$ for various $\lambda$ values. Here, $t_{so}=0.5$, $M=-4$, and $\phi=0$.

### v3

@@ Abstract
Topological Anderson insulators (TAIs) provide a mechanism for topological phase transitions in disordered systems and have implications for quantum material design. In this work, we investigate the emergence of multiple TAIs in a one-dimensional spin-orbit coupled (SOC) chain subject to Fibonacci modulation, which transforms a trivial band structure into a sequence of topologically nontrivial phases. This behavior is characterized by the appearance of zero-energy modes and changes in the $Z_{2}$ topological quantum number. As the SOC amplitude decreases, the number of TAI phases increases, a feature that is closely related to the fractal structure of the energy spectrum induced by Fibonacci modulation. In contrast to conventional TAI phases with fully localized eigenstates, the wave functions in the Fibonacci-modulated TAI phases display multifractal properties. This model can be experimentally realized using a Bose-Einstein condensate in a momentum-space lattice, where its topological transitions and multifractal features can be explored through quench dynamics.
Introduction
Topological states of matter have been extensively studied in various systems, including condensed-matter materials MZH2010 ; zhangSC2011 ; DasT2016 ; Ryu2016 ; Schindler2024 ; SlagerRJ202401 ; SlagerRJ202402 , ultracold atoms Spielman2019 ; Bhattacharya2023 , superconducting circuits Lehnert2014 ; Tan2019 ; Sun2019 ; Tan2021 , electronic circuits zhangxd2021 , photonic lattices Lul2014 ; Ozawa2019 ; Vicencio2024 ; Qu2024 ; Rechtsman2024 , and mechanical systems Huber2016 . Topological insulators are characterized by gapless edge states that are immune to backscattering from weak disorders Raghu2008 ; Wangz2008 ; Wangz2009 , but these states typically cannot survive strong disorders due to Anderson localization Chong2017 ; Zhangs2017 . A reverse transition has been observed: disorder can induce the emergence of protected edge states and quantized transport even in a trivial band structure. This disorder-induced topological phase, known as the topological Anderson insulator (TAI) Shen2009 ; Groth2009 ; Wang2011 ; Shen2012 ; Titum2015 ; Orth2016 ; Jiang2019 ; Huang2022 ; Cui2022 ; Sun2024 ; Luzp2025 , has been theoretically predicted in systems such as the two-dimensional (2D) HgTe/CdTe quantum well Shen2009 ; Groth2009 ; Xie2009 and the 1D Su-Schrieffer-Heeger chain Schmiedt2014 , and experimentally realized in a variety of engineered systems, including 1D cold atomic wires Gadway2018 , 2D photonic waveguide arrays Szameit2018 ; LiuGG2020 , and photonic quantum walk XueP2022 . Both uncorrelated random Gadway2018 ; Dunlap1990 ; ZhangDW202202 ; Nava2023 ; Cinnirella2024 ; KangDW2024 ; GhoshAK2024 and quasiperiodic disorders in the form of a cosine modulation incommensurate with lattice spacing ZhuSL2021 ; ZhangDW2022 ; LuZP2022 ; Luzp2025 ; Padhan2024 ; XueQK2024 can give rise to the TAI phase. Expanding the range of disorder types that can induce the TAI phase, and exploring the associated topological and localization effects, remains a critical focus of ongoing research.
Quasicrystals provide platforms for studying exotic localization phenomena and topologically nontrivial behavior in physics Ouyang2024 ; DaiQ2023 ; YaoH2020 ; Modugno2014 ; Garreau2008 ; Rai2021 ; Haas2024 ; Kobialka2024 ; Sandberg2024 . Unlike randomly disordered systems, quasicrystals possess long-range order without periodicity Sbroscia2020 ; Jagannathan2021 . Recent advances in quasicrystal growth and synthetic engineering—achieved with atomic and photonic precision—have enabled the realization of quasiperiodic structures, among which the Fibonacci quasicrystal stands out as a prominent example. In a conventional Fibonacci-modulated chain, several features emerge that are characteristic of quasicrystalline order. The energy spectrum displays a self-similar, fractal structure known as a ”Cantor set”, which is neither entirely continuous nor discrete. This complex spectrum is accompanied by a highly nontrivial spatial structure of the eigenstates: single-particle wave functions are typically critical, meaning they are neither fully localized nor completely extended 1983 ; Zilberberg2012 ; Liutong2025 ; Verbin2013 ; Mace2016 ; Mace2017 ; Bian2021 ; Jagannathan2021 . Such critical states arise directly from the underlying quasiperiodic order and result in multifractal properties for both the spectrum and the eigenfunctions. Furthermore, the system lacks translational symmetry but retains long-range order, giving rise to unique transport and localization phenomena that are distinct from those found in periodic or purely disordered systems. These features make the Fibonacci chain a paradigmatic platform for exploring the interplay between order, disorder, and topology in low-dimensional quantum systems Rai2021 ; Haas2024 ; Kobialka2024 ; Sandberg2024 . Synthetic Fibonacci chains have been realized in various platforms, including cold atomsSingh2015 ; Weld2024 , photonics Verbin2013 ; Kraus2012 ; Negro2003 ; Verbin2015 , polaritonics Tanese2014 ; Goblot2020 , and dielectric chains or circuits Reisner2023 ; Franca2024 , where the wave functions and topological properties of Fibonacci chains have been directly measured and investigated experimentally Dareau2017 ; Rai2021 ; LiuF2018 ; LiuF2019 ; XuDH2020 .
In this work, we report the emergence of multiple TAIs induced by Fibonacci modulation in a 1D spin-orbit coupled (SOC) chain, starting from a trivial band structure. This phenomenon is confirmed by the appearance of nontrivial zero modes, along with changes in the $\mathcal{Z}_{2}$ topological quantum number. The number of TAI phases increases as the SOC amplitude decreases. The fractal structure of the spectrum plays a crucial role in governing the emergence of these multiple TAI phases. Specifically, the self-similar splitting of the spectrum induced by the Fibonacci modulation facilitates the formation of multiple topologically nontrivial regions, which evolve as the SOC amplitude is varied. In contrast to conventional TAI phases, which exhibit fully localized eigenstates, the wave functions in the Fibonacci-modulated TAI phases display multifractal characteristics. Furthermore, our model can be experimentally realized in a Bose-Einstein condensate along the momentum lattice, where its topological transitions and multifractal behavior can be probed through quench dynamics.
Results and discussion
Model and Hamiltonian
@@ Introduction
We investigate a 1D SOC atomic chain subjected to a Fibonacci-modulated on-site potential. This system is described by the Hamiltonian
@@
where $\Psi_{i}^{\dagger}=(c^{\dagger}_{i\uparrow},c^{\dagger}_{i\downarrow})$, and $c^{\dagger}_{i\beta}$ creates a particle with spin $\beta=\{\uparrow,\downarrow\}$ at site $i$. The hopping matrix is given by $\mathcal{R}_{ij}=-t_{0}\sigma_{z}\pm it_{so}\sigma_{y}$, corresponding to hopping along the $\pm\hat{x}$ direction, respectively. Here, the diagonal elements of $\mathcal{R}$ describe spin-conserved hopping with amplitude $t_{0}$ (set as the energy unit, $t_{0}=1$), while the off-diagonal terms represent the spin-flip hopping between nearest neighbors with amplitude $t_{so}$. The second term in the Hamiltonian accounts for the on-site Fibonacci modulation Jagannathan2021 , expressed as
$\mathcal{U}_{i}=\left[\lambda\mathrm{sgn}\left(\cos(2\pi\alpha i+\phi)-\cos(\pi\alpha)\right)+M\right]\sigma_{z},$ (2)
where $M$ is a uniform Zeeman potential, $\lambda$ denotes the amplitude of quasiperiodic modulation, $\textup{sgn}[\cdots]$ is the sign function, and $\phi\in[0,2\pi)$ is an arbitrary phase used to generate different modulation configurations. A shift in the phase $\phi$ is equivalent to a translation in the site index and can therefore be eliminated by redefining the origin of the position. The modulation frequency $\alpha$ is chosen as an irrational Diophantine number, which can be constructed from a generalized $\kappa$-Fibonacci sequence defined by $F_{\nu+1}=F_{\nu-1}+\kappa F_{\nu}$ with $F_{0}=F_{1}=1$ Kohmoto1983 . The corresponding irrational number is given by $\alpha=\lim_{\nu\to\infty}F_{\nu-1}/F_{\nu}$, where $\kappa=1,2,3,\cdots$, yielding the so-called metallic mean family. For example, the golden mean ($\alpha_{g}=(\sqrt{5}-1)/2$) corresponds to $\kappa=1$, the silver mean ($\alpha_{s}=(\sqrt{2}-1)$) to $\kappa=2$, and the bronze mean ($\alpha_{b}=(\sqrt{13}-3)/2$) to $\kappa=3$, and so on. In our study, the system size is set as $L=F_{\nu}$ and the modulation frequency is approximated by the rational number $\alpha=F_{\nu-1}/F_{\nu}$. This generalized $\kappa$-Fibonacci sequence of quasiperiodic modulations can be constructed using various methods, including the substitution method, the cut-and-project method Ouyang2024 ; Jagannathan2021 , and the characteristic function method (see Supplementary Note 1 for details). Here, Eq. (2) corresponds to the characteristic function method.
@@
The Hamiltonian $H$ respecting the chiral symmetry defined by $\sigma_{x}$ in spin space governs a $\kappa$-Fibonacci-modulated SOC chain. Its counterpart with a constant uniform Zeeman potential realizes the 1D AIII class topological insulator LiuXJ2013 . In the crystalline case, where $\mathcal{U}_{i}=M\sigma_{z}$, the clean SOC model splits into two bands with eigenvalues $E_{\pm}=\pm\sqrt{\left(M-2\cos k\right)^{2}+4t_{so}^{2}\sin^{2}k}$, where $k$ is the wave vector. For $|M|>2$, regardless of whether $t_{so}\neq 0$, a topological phase transition occurs, indicated by the closing of the energy gap at $\left|M\right|=2$. This transition is further characterized by the vanishing of zero-energy edge modes and the disappearance of nontrivial topological invariants. In this work, we primarily focus on the case of $\kappa=1$, investigating the emergence of multiple TAI phases and multifractal features induced by the Fibonacci modulation. We also examine the cases with $\kappa=2$ and $3$ (see Supplementary Note 4 for details), where the presence of multiple TAI phases is similarly observed.
TAI and multiple TAIs
@@
Topologically protected edge states and quantized topological charges emerge in the context of TAI when sufficient disorder or incommensurate modulation is introduced to a trivial band structure. We employ the scattering-matrix method to compute the $\mathcal{Z}_{2}$ topological quantum number $Q$ in our case (see Methods), where $Q=1$ (Q=1Q=-1) corresponds to a topologically trivial (nontrival) phase. Figure 1(a) shows the disorder-averaged topological phase diagram with $t_{so}=0.5$ and $\kappa=1$ as a function of $M$ and $\lambda$, obtained numerically averaging the disorder-averaged $\mathcal{Z}_{2}$ number $\overline{Q}$ over $N_{c}$ disorder realizations for different $\phi\in[0,2\pi)$. The phase boundaries can be determined by the condition:
@@
where $T$ is the total transfer matrix defined in a new basis $\tilde{\Psi}_{i}^{\dagger}=(c_{i\uparrow}^{\dagger}+c_{i\downarrow}^{\dagger},c_{i\uparrow}^{\dagger}-c_{i\downarrow}^{\dagger})$ for the zero modes (see Supplementary Note 2 for details). In the absence of $\lambda$, the system exhibits a topological phase transition at $|M|=2$, corresponding to a jump in $\overline{Q}$ from 1-1 to $1$ as $|M|$ increases. When the Fibonacci modulation ($\kappa=1$) is introduced to the nontrivial band structure with $|M|<2$, the topologically nontrivial phase remains robust against weak $\lambda$, but eventually transitions to a trivial one for sufficiently strong modulation. For $M>2$, the system undergoes a topological phase transition from trivial to nontrivial to trivial as $\lambda$ increases, indicating the emergence of the TAI. We further present two disorder-averaged energies, $\overline{E}_{L}$ and $\overline{E}_{L+1}$, at the center of the spectrum, as well as $\overline{Q}$ as a function of $\lambda$ under open boundary condition (OBC) with $\kappa=1$, $M=3$ and $t_{so}=0.5$ in Fig. 1(b). When $\lambda$ lies in the range $[2.51,2.94]$, the topological number $\overline{Q}$ jumps from $1$ to 1-1, accompanied by the emergence of the zero edge modes. For $\lambda>2.94$, the topologically nontrivial phase vanishes. Our numerical results suggest that the Fibonacci modulation can induce the TAI.
Figure 1: Topological transitions for different modulation and SOC strengths. (a) The disorder-averaged topological phase diagram with the spin-flip hopping strength $t_{so}=0.5$ as a function of the uniform Zeeman potential $M$ and the quasiperiodic modulation amplitude $\lambda$. The blue dashed lines correspond to the topological phase boundaries determined by Eq.(3). The black (red) dashed line corresponds to the line of $M=3$ (M=4)(M=-4). Two disorder-averaged energies $\overline{E}_{L}$ and $\overline{E}_{L+1}$ at the center of the spectrum and the disorder-averaged topological number $\overline{Q}$ as a function of $\lambda$ under OBC with $t_{so}=0.5$ for (b) $M=3$ and (c) M=4M=-4. (d) The number of times the TAI phases emerge, denoted as $\overline{g}$, as a function of $t_{so}$ with M=4M=-4. Here, $\kappa=1$, and all the data are averaged by $N_{c}=50$ disorder realizations.
Figure 2: Energy spectra of the system with Fibonacci modulation for different SOC strengths. (a) Self-similarities of energy spectrum as a function of the quasiperiodic modulation amplitude $\lambda$ with the spin-flip hopping strength $t_{so}=0$. The insets in (a) show magnified views of the regions within the red boxes. Two central energy levels $E_{L}$ and $E_{L+1}$ as a function of $\lambda$ with (b) $t_{so}=0.25$, (c) $0.5$, and (d) $0.9$. The inset in (b) shows a magnified view of the original image. Here, the uniform Zeeman potential M=4M=-4, $\kappa=1$, $\phi=0$, and $L=610$.
@@
As seen in Fig. 1(a) for M<2M<-2, the system undergoes multiple distinct transitions into the TAI phase, eventually returning to a trivial phase as $\lambda$ increases further. Figure 1(c) shows the disorder-averaged energies of the two center modes, $\overline{E}_{L}$ and $\overline{E}_{L+1}$, as well as $\overline{Q}$ as a function of $\lambda$ with $\kappa=1$, M=4M=-4 and $t_{so}=0.5$. It can be observed that, as the $\mathcal{Z}_{2}$ number $\overline{Q}$ transitions to a nontrivial value, corresponding zero modes emerge in the ranges $\lambda\in(2.87,3.22)\cup(3.67,3.94)\cup(4.66,4.88)$. The system undergoes three transitions from topologically trivial phases to nontrivial ones, indicating the emergence of the multiple TAI phases. For a given M<2M<-2, the number of times TAI phases emerge, denoted by $g$, strongly depend on $t_{so}$. As shown in Fig. 1(d) with $\kappa=1$ and M=4M=-4, $\overline{g}$ gradually increases as $t_{so}$ decreases, for $N_{c}=50$ disorder realizations.
@@
To uncover the underlying mechanism of this phenomenon, we analyze the energy spectrum of two decoupled Fibonacci-modulated chains ($\kappa=1$) without the SOC term ($t_{so}=0$) as described in Eq. (1), shown in Fig. 2(a) for $\phi=0$. For finite $\lambda$, the spectrum exhibits eight primary bands. Within the range $\lambda\in(2.76,4.95)$, the middle six bands overlap, forming three clusters of band-crossing regions separated by two prominent band gaps near zero energy. Remarkably, each cluster undergoes successive splitting into smaller sub-clusters, creating a self-similar fractal structure that persists in the thermodynamic limit. This fractal nature manifests in the emergence of progressively smaller band gaps $\Delta_{g}$ at each level of splitting. When the SOC strength $t_{so}$ is introduced, these band gaps comparable in size to $t_{so}$ begin to close and reopen, signaling the formation of nontrivial zero-energy modes, as illustrated in Fig. 2(b). This process unveils an infinite hierarchy of TAI phases in the small $t_{so}$ regime and in the thermodynamic limit, effectively creating a topologically fractal structure. As $t_{so}$ increases further, neighboring topologically nontrivial regions merge, reducing the number of distinct TAI phases. This transition, depicted in Figs. 2(c) and 2(d), signifies a transformation from a fractal topology to a more discrete structure, highlighting how the interplay between SOC and Fibonacci modulation reshapes the landscape of topological phases.
In Supplementary Note 3, we present an off-diagonal Fibonacci-modulated SOC chain ($\kappa=1$) and uncover the emergence of multiple TAI phases, whose occurrence strongly depends on $t_{so}$. The localization properties exhibit multifractal behavior, regardless of whether the system resides in a TAI phase or not for the off-diagonal Fibonacci-modulated SOC chain. These findings demonstrate that both diagonal and off-diagonal Fibonacci modulation in SOC chains can induce multiple TAI phases, offering insights into quasiperiodic topological systems. To study the modulation contribution to the emergence of TAIs, we investigate the cases with the modulation frequency chosen as other metallic mean numbers ($\kappa=2$ and $3$) in Supplementary Note 4. Our results show that, similar to the golden mean case, both the silver mean and bronze mean modulations lead to multiple reentrant TAI phases as the disorder strength increases. The topological phase diagrams for these cases reveal a sequence of transitions between trivial and topological phases, with the number of TAI regions and their evolution strongly affected by the modulation frequency and SOC strength. The energy spectra without SOC exhibit self-similar fractal structures. Introducing finite SOC causes band gaps to close and reopen, resulting in the emergence and eventual merging of TAI regions. These findings confirm that the appearance of multiple reentrant TAI phases is a universal and robust phenomenon for various metallic mean modulations, with the modulation frequency playing a key role in shaping the complexity and evolution of the topological phase diagram.
Localization properties
Figure 3: Density distributions in both real and momentum spaces. Density distributions of selected eigenstates for various values of the quasiperiodic modulation amplitude $\lambda$ in both real (a1)-(a6) and momentum (b1)-(b6) spaces with the spin-flip hopping strength $t_{so}=0.5$, the uniform Zeeman potential M=4M=-4, $\kappa=1$, $\phi=0$, and $L=610$. From left to right, the columns correspond to $\lambda=2.0$, $3.0$, $3.5$, $3.8$, $4.2$, and $4.8$, where the system is localized in different topological regimes.